### v1

# Angular Momentum Drain: Despinning Embedded Planetesimals

Abstract
Young and forming planetesimals experience impacts from particles present in a protostellar disk. Using crater scaling laws, we integrate ejecta distributions for oblique impacts. For impacts at 10 to 65 m/s, expected for impacts associated with a disk wind, we estimate the erosion rate and torque exerted on the planetesimal. We find that the mechanism for angular momentum drain proposed by Dobrovolskis and Burns (1984) for asteroids could operate in the low velocity regime of a disk wind. Though spin-down associated with impacts can facilitate planetesimal collapse, we find that the process is inefficient. We find that angular momentum drain via impacts operates in the gravitational focusing regime, though even less efficiently than for lower mass planetesimals. The angular momentum transfer is most effective when the wind speed is low, the projectile density is high compared to the bulk planetesimal density, and the planetesimal is composed of low-strength material. Due to its inefficiency, we find that angular momentum drain due to impacts within a pebble cloud does not by itself facilitate collapse of single planetesimals.
1 Introduction
The spin rates of most asteroids have been altered by impacts (Farinella et al., 1992; Mao and McKinnon, 2020). Because high velocity ejecta preferentially escapes along the direction of rotation, impacts cumulatively cause asteroids and other minor bodies to spin down (Dobrovolskis and Burns, 1984; Cellino et al., 1990; Takeda and Ohtsuki, 2009; Ševeček et al., 2019). Additional mechanisms for minor body spin evolution include sublimation and outgassing (Mueller and Ferrin, 1996; Mottola et al., 2020), the YORP effect (Bottke et al., 2006) and tidal encounters (Scheeres et al., 2004). In this study we focus on the mechanism for angular momentum drain by impacts proposed by Dobrovolskis and Burns (1984), but apply it in the context of recently formed planetesimals during the epoch when they are embedded in a circumstellar disk.
Planetesimals are thought to form in particle-rich clumps that coalesce within a protoplanetary disk during an era of streaming instability (Johansen et al., 2007; Li et al., 2019; Carrera et al., 2021) or within turbulent vortices (Cuzzi et al., 2008). For variations or alternatives to these scenarios, see the review by Simon et al. (2022). Planetesimal collapse within a cloud of pebbles is expected to be inefficient, with only a fraction (between 50% to 100%) of a pebble cloud mass incorporated into bound objects (Robinson et al., 2020; Nesvorný et al., 2021; Polak and Klahr, 2023). During the final states of planetesimal collapse, the angular momentum of the pebble cloud can generate a rapidly spinning disk or a bound binary (Nesvorný et al., 2019; Robinson et al., 2020; Nesvorný et al., 2021; Polak and Klahr, 2023). A newly formed planetesimal can experience accretion, erosion, and transport of particles across its surface while embedded in the protoplanetary disk (Rozner et al., 2020; Demirci et al., 2020; Quillen et al., 2024).
For a clump of mass $M$ to be self-gravitating, its self-gravity should exceed the tidal force from the host star. A gas or particle cloud would collapse if its mean density $\rho_{c}\gtrsim\rho_{H*}$ where the modified Hill density is
$\rho_{H*}$ $=\frac{1}{\pi}\left(\frac{3}{2}\right)^{5}\frac{M_{*}}{a^{3}},$ (1)
$=1.4\times 10^{-3}{\rm\ kg~{}m}^{-3}\left(\frac{M_{*}}{M_{\odot}}\right)\left(\frac{a}{1\ {\rm AU}}\right)^{-3}$
(Paczyński 1971, but also following equation 2 by Polak and Klahr 2023). Here $M_{*}$ is the mass of the star and $a$ is the orbital semi-major axis about the star. At an orbital semi-major axis of $a=45$ AU with $M_{*}=M_{\odot}$, the modified Hill density is $\rho_{H*}=1.54\times 10^{-8}{\rm\ kg~{}m}^{-3}$. This is approximately $10^{11}$ times smaller than the density of Kuiper Belt Objects, ranging from $235\ {\rm kg~{}m^{-3}}$ to $2500\ {\rm kg~{}m^{-3}}$ (Brown, 2012; Bierson and Nimmo, 2019; Keane et al., 2022).
If a self-gravitating pebble cloud has some spin angular momentum, during collapse it would spin up (e.g., Nesvorný et al. 2021). If the angular momentum of the cloud exceeds that of a critically rotating Jacobi ellipsoid of density about 1000 kg m-3, then angular momentum must be shed for the object to collapse into a single bound object.
In the context of the formation of Transneptunian binaries, mechanisms for removing angular momentum include the ejection of a third body (Noll et al., 2008), gravitational interactions with a distribution of small bodies residing in the disk (Goldreich et al., 2002; Funato et al., 2004) or collisions (Weidenschilling, 2002). The simulations by Nesvorný et al. (2021) illustrate that the particles within the pebble cloud itself can serve as a reservoir for angular momentum, and that collisions between particles within the cloud can facilitate dissipation and angular momentum transfer. Binaries, contact binaries, and disk shaped objects may be a consequence of the angular momentum of a pebble cloud prior to its collapse (Nesvorný et al., 2021; Polak and Klahr, 2023).
We focus on the connection between impacts from particles in a pebble cloud and the spin angular momentum of a collapsing body. As impacts can spin down an asteroid (Dobrovolskis and Burns, 1984), impacts could aid in collapse of a pebble cloud, and the host pebble cloud would serve as a reservoir or sink for angular momentum. While planetesimals are embedded in a disk, low velocity impacts could be frequent (Quillen et al., 2024). While impacts in the asteroid belt tend to be hypervelocity at a few km/s (Bottke et al., 1994) and can generate a spray of ejected particles (Lauretta et al., 2019), impact velocities on a planetesimal within a pebble cloud would be substantially lower, between a few to tens of m/s (Johansen et al., 2007; McKinnon et al., 2020; Quillen et al., 2024).
With a high velocity impact on a small body, most ejecta escapes. But at lower impact velocities, the fraction of ejecta that escapes is more sensitive to body rotation, impact location, impact and ejecta angles, and the planetesimal’s escape velocity (Hirata et al., 2021; Hyodo and Genda, 2021; Quillen et al., 2024). Using an ejecta velocity distribution, we can determine the mass of ejecta that is above the escape velocity and so escapes the planetesimal. The work by Dobrovolskis and Burns (1984) dealt with high-velocity, erosional impacts that reduce the spin-rate of the asteroid. We extend this work to a low velocity regime to see how impacts affect the mass loss and angular momentum of a planetesimal that is embedded in a protostellar disk.
The integrations by Dobrovolskis and Burns (1984) neglected the sensitivity of ejecta mass to azimuthal angle for oblique impacts, however recent studies (Raducan et al., 2022; Luo et al., 2022; Quillen and Doran, 2024) have begun to characterize the azimuthal dependence of oblique impact ejecta distributions. In this work, we use recent estimates for the azimuthal sensitivity of ejecta to impact angle to determine which ejecta escape. Dobrovolskis and Burns (1984) focused on the role of impacts on asteroids. In this setting the projectiles originate from different directions. The changes in angular momentum caused by impacts were modeled with a random walk (Dobrovolskis and Burns, 1984). In contrast, the spin axis of a planetesimal that is embedded within a circumstellar disk could be related to the direction of a headwind containing particle projectiles (Nesvorný et al., 2021; Quillen et al., 2024). Consequently in this study we consider impacts from a uniform direction that is perpendicular to the planetesimal spin axis.
To form planetesimals within a protostellar disk, dust particles and ices coagulate to form low density and low-strength aggregates, which are often called pebbles (Lorek et al., 2018; O’Rourke et al., 2020). Millimeter sized aggregates, that later form into planetesimals, initially would be porous, with volume filling factor well below a percent Lorek et al. (2018). An aggregate growing via cohesive or sticky collisions reaches a size known as the bouncing barrier, where collisions bounce rather than stick, compressing the aggregate to a volume filling factor of 1 - 10% (Lorek et al., 2018). Icy and dusty aggregates are likely porous or fluffy when the disk undergoes streaming instability (Lorek et al., 2018). Because of the large range of possible densities for pebbles in a primordial disk, in this study we also vary the ratio of projectile to planetesimal density.
1.1 Outline
In Section 2 we discuss reference frames and planetesimal collisions with particles that originate in a headwind. We use the point-source approximation ejecta scaling model developed for normal impacts (Housen and Holsapple, 2011) that has been extended to be sensitive to impact angle as well as ejecta azimuthal angle (Raducan et al., 2022; Quillen and Doran, 2024). We describe how we estimate the angular momentum transfer and fraction of ejecta that escape from single impacts.
In Section 3 we apply the framework for single impacts to a distribution of collisions across the planetesimal surface. These collisions are oriented on a Cartesian grid overlaid onto the cross-sectional area of the planetesimal in the inertial reference frame. We use the Transneptunian object Arrokoth as a fiducial setting for our integrations and highlight the importance of gravitational focusing.
In Section 4, we vary planetesimal and projectile parameters to explore the sensitivity of the angular momentum transfer efficiency and mass loss ratio to physical properties such as planetesimal and projectile density, planetesimal strength, planetesimal spin rate, and projectile velocity, and spatial projectile velocity gradient.
2 Impacts on an Embedded Planetesimal
We consider a planetesimal that is bombarded by particles that reside in a circumstellar disk. We assume that, due to a difference between the velocity of the planetesimal and that of gas and small particles in the disk, impacts originate from a preferred direction, as shown in Figure 1. A planetesimal embedded in the disk feels a wind known as a ‘headwind’ that arises because the gas in the disk feels pressure, giving the gas a mean tangential circular velocity below that of a circular Keplerian orbit (Whipple, 1972). We refer to the particle flow as particles within a headwind, but impacts from particles could also occur within a vortex clump (Cuzzi et al., 2008) or a pebble cloud clump that has formed during streaming instability (Johansen et al., 2007).
2.1 Planetesimal Properties
Following Dobrovolskis and Burns (1984), we take the planetesimal to be spherical with uniform escape velocity,
$u_{esc}=\sqrt{2GM_{a}/R_{a}}$ (2)
where $M_{a}$ is the planetesimal mass, $R_{a}$ is the planetesimal radius, and $G$ is the gravitational constant. We assume solid body rotation and an isotropic and homogeneous body.
Because planetesimals can have a high spin rate (Johansen and Lacerda, 2010; Robinson et al., 2020; Nesvorný et al., 2021), we describe body spin in terms of its maximum rotation rate where a particle on the equator would leave the surface due to centrifugal acceleration. We put spin in units of the centrifugal break-up value for a sphere,
$\Omega_{\rm break-up}$ $=\sqrt{\frac{4\pi G\rho_{a}}{3}},$ (3)
where $\rho_{a}$ is the planetesimal’s bulk density.
Figure 1: An image indicating the inertial reference frame coordinate system. The direction of planetesimal rotation is shown with respect to the headwind. Dashed lines show trajectories of headwind particles that collide with the planetesimal. We assume that the planetesimal spin axis is perpendicular to the circumstellar disk plane.
The rotation axis of the planetesimal is assumed to be perpendicular to the incident wind direction, as shown in Figure 1. This assumption is motivated by simulations of pebble accretion which suggest that planetesimals form with rotation axis prograde with respect to the disk in which they are embedded (Johansen and Lacerda, 2010). Figure 1 illustrates a Cartesian inertial coordinate system, with wind particles originating from the $+x$ direction and the planetesimal spin axis pointing toward the $+z$ direction.
2.2 Inertial and Rotation Frames
The impact velocity for a collision with a headwind particle is 10 - 65 m/s (e.g., Quillen et al. 2024) and is low compared to most impacts that currently take place in the inner solar system, which are at velocities of a few km/s (Bottke et al., 1994). A 10 km planetesimal of bulk density $\rho_{a}=1000$ kg m-3 spinning near breakup has rotational velocity of about 5 m/s. The ratio of rotational velocity to impact velocity and the ratio of rotational velocity to the velocity of ejecta are higher in our setting than in the context of km/s impacts on asteroids (studied by Dobrovolskis and Burns 1984). As a consequence we do not neglect the rotational velocity when computing impact velocities in the frame rotating with the planetesimal surface. Figure 2 illustrates how the impact velocity, as seen in a frame rotating with the surface, depends upon the impact location and the direction of rotation.
We denote the location of an impact on the planetesimal surface with the vector ${\bf r}$, originating from the planetesimal’s center of mass. The velocity of the surface at ${\bf r}$ is
${\bf u}_{rot}={\bm{\Omega}}_{a}\times{\bf r}.$ (4)
where ${\bm{\Omega}}_{a}$ is the planetesimal spin vector, which we take to be aligned with $\hat{z}$. On a spherical planetesimal ${\bf r}=R_{a}\hat{\bf n}.$
2.3 The Impact Velocity and Impact Angle in the Rotating Frame
Figure 2: We illustrate how rotation affects impact angle and velocity in as seen in frame rotating with the surface. Black shows rotation. Blue shows wind speed. Green shows wind particle velocity in the frame rotating with the planetesimal. From the perspective of a viewer on the surface on the leading side, rotation causes the wind particle to impact the surface at higher velocity and at lower (closer to grazing) impact angle.
Figure 3: In the frame rotating with the surface, we illustrate variables relevant for describing crater, projectile, and ejecta. The variables $x,\zeta$ are coordinates measured with origin at the site of impact. The normal unit vector $\hat{\bf{n}}$, the down-range unit vector $\hat{\bf{d}}$, and $\hat{\bf{b}}$ are orthogonal vectors defining a local coordinate system. The crater radius $R_{cr}$ is the radius from the site of impact where the crater profile crosses the surface level (before impact) and it can be a function of impact angle $\theta_{imp}$ and azimuthal angle $\zeta$. The ejecta velocity is $v_{ej}$.
The ejecta velocity distribution depends on both impact velocity and direction in the frame rotating with the surface at the point of impact. In the rotating frame moving with the surface, the projectile velocity is
${\bf v}_{pj}={\bf u}_{pj}-{\bf u}_{rot}$ (5)
where ${\bf u}_{pj}$ is the projectile velocity in the inertial frame, as shown in Figure 1. The rotation velocity ${\bf u}_{rot}$ is calculated using Equation 4 for a given position vector ${\bf r}$.
We define a unit vector $\hat{\bf n}$ as vector normal to the surface and pointing away from the surface at the site of impact. We define a downrange vector $\hat{\bf d}$ as a unit vector tangential to the surface pointing downrange (or away) from the direction from which the projectile originated. The unit normal and downrange vectors are
$\hat{\bf{n}}$ $=\frac{{\bf r}}{r}$ (6)
$\hat{\bf d}$ $=\frac{{\bf v}_{pj}-({\bf v}_{pj}\cdot\hat{\bf n})\hat{\bf n}}{|{\bf v}_{pj}-({\bf v}_{pj}\cdot\hat{\bf n})\hat{\bf n}|}$ (7)
where ${\bf v}_{pj}$ is the projectile velocity vector in the frame rotating with the surface. The unit vector $\hat{\bf d}$ is computed by projecting the incoming projectile velocity into the plane tangent to the surface at the point of impact. It is convenient to have a set of 3 orthogonal unit vectors at the site of impact. A third unit vector, $\hat{\bf b}$, called the binormal vector, is perpendicular to $\hat{\bf n}$ and $\hat{\bf d}$, and is
$\hat{{\bf b}}=\hat{{\bf n}}\times\hat{{\bf d}}.$ (8)
The impact angle, $\theta_{imp}$ is defined as the angle between the projectile velocity (in the rotating frame) and the plane tangent to the surface
$\theta_{imp}=\arcsin{\left(\frac{-{\bf v}_{pj}\cdot\hat{\bf n}}{|{\bf v}_{pj}|}\right)}.$ (9)
The minus sign arises because we have defined the normal direction $\hat{\bf n}$ pointing outward from the surface. The impact angle is low for grazing impacts and equal to $\pi/2$ for a normal impact. Figure 3 illustrates the different directions and angles for an impact in the rotating frame.
2.4 Crater Scaling and Ejecta Distributions
The momentum transfer parameter and the angular momentum imparted to a planetesimal by an impact depends upon the integrated properties of the ejecta (e.g., Dobrovolskis and Burns 1984; Takeda and Ohtsuki 2009; Raducan et al. 2022). We first discuss the velocity and mass distributions of ejecta from normal impacts and then consider more general ejecta distributions from oblique impacts.
When computing the properties of ejecta, we assume that planetesimal mass exceeds projectile mass, $M_{a}\gg m_{pj}$, and planetesimal mass also exceeds total ejecta mass; $M_{a}\gg M_{ej}$. We neglect the possibility of projectile ricochet (e.g., Wright et al. 2020) and projectile spin (e.g., Carvalho et al. 2023a, b, 2024).
Normal impacts are often described in terms of three dimensionless parameters (Housen and Holsapple, 2011);
$\pi_{2}$ $=\frac{g_{a}a_{pj}}{v_{pj}^{2}}$ (10)
$\pi_{3}$ $=\frac{Y_{a}}{\rho_{a}v_{pj}^{2}}$ (11)
$\pi_{4}$ $=\frac{\rho_{a}}{\rho_{pj}}$ (12)
where $g_{a}$ is the surface gravitational acceleration on the planetesimal surface, the projectile radius is $a_{pj}$, and the density of planetesimal and projectile are $\rho_{a}$ and $\rho_{pj}$, respectively. Impacts from headwind particles onto a planetesimal are likely to be in the strength regime (Quillen et al. 2024 following Scheeres et al. 2010) where crater and ejecta properties only depend upon $\pi_{3}$ and $\pi_{4}$.
We adopt the ejecta scaling laws for normal impacts by Housen and Holsapple (2011), which were derived assuming a point source approximation. The ejection velocity is a function of distance $x$ from the impact site where the ejecta crosses the surface level plane (as it was prior to impact), as illustrated in Figure 3. The ejecta velocity scaling law is
$\frac{v_{ej,N}(x,\pi_{4})}{v_{pj}}$ $=C_{1,N}\left(\frac{x}{a_{pj}}\right)^{-\frac{1}{\mu_{N}}}\pi_{4}^{-\frac{\nu}{\mu_{N}}}\left(1-\frac{x}{n_{2}R_{cr,N}}\right)^{p}.$ (13)
The exponents $\mu_{N},\nu$ and coefficients $C_{1,N}$ depend upon substrate material properties. We have added an $N$ subscript, denoting normal impact, to parameters that we vary below when discussing oblique impacts.
Housen and Holsapple (2011) find that the coefficients $n_{1}\sim 1.2$, $n_{2}\sim 1$, and $p\sim 0.3$. Equation 13 holds for $n_{1}a_{pj}<x<n_{2}R_{cr,N}$, where $R_{cr,N}$ is the crater radius for a normal impact. The crater radius is the radius at which the crater profile crosses the surface level plane, as shown in Figure 3. In the strength regime, the crater radius for a normal impact is
$R_{cr,N}(\pi_{3},\pi_{4})=a_{pj}\left(\frac{4\pi}{3}\right)^{\frac{1}{3}}H_{2}\pi_{3}^{-\frac{\mu_{N}}{2}}\pi_{4}^{-\nu},$ (14)
with coefficient $H_{2}\sim 0.4$ that also depends upon substrate material properties (Housen and Holsapple, 2011).
The ejecta mass integrated between $n_{1}a_{pj}<x<n_{2}R_{cr,N}$ from the site of impact is
$\frac{M(<x,\pi_{4})}{m_{pj}}=\frac{3k_{N}}{4\pi}\pi_{4}\left(\left(\frac{x}{a_{pj}}\right)^{3}-n_{1}^{3}\right)$ (15)
and with coefficient $k_{N}\sim 0.3$ (Housen and Holsapple, 2011).
2.4.1 Ejecta distributions for oblique impacts
Ejecta velocity, ejecta angle, and mass distributions for oblique impacts are sensitive to both impact angle, $\theta_{imp}$, and azimuthal angle within the crater, $\zeta$ (Anderson et al., 2003; Raducan et al., 2022; Quillen and Doran, 2024). We take azimuthal angle $\zeta$ to be the angle in the surface plane (as seen from above the crater in the plane spanned by unit vectors $\hat{\bf d}$ and $\hat{\bf b}$), with direction $\zeta=180^{\circ}$ downrange, as adopted by Anderson et al. (2003); Raducan et al. (2022).
Following Raducan et al. (2022); Quillen and Doran (2024) we modify equations 13, 14, and 15 to be sensitive to both impact angle and azimuthal angle, giving functions for crater radius $R_{cr}$, exponent $\mu$, and coefficients $C_{1},k$ (dropping the $N$ subscript) that are consistent with normal impact scaling
$\mu(\zeta,\theta_{imp}=\pi/2)$ $=\mu_{N}$ (16)
$C_{1}(\zeta,\theta_{imp}=\pi/2)$ $=C_{1,N}$ (17)
$R_{cr}(\zeta,\theta_{imp}=\pi/2)$ $=R_{cr,N}$ (18)
[…]
We adopt the functions estimated by Raducan et al. (2022) from measurements of ejecta from 3D iSALE oblique impact simulations;
$\mu(\zeta,\theta_{imp})$ $=\mu_{N}\left(1+\frac{1}{2}\cos\zeta\cos\theta_{imp}\right)$ (20)
$C_{1}(\zeta,\theta_{imp})$ $=C_{1,N}\exp(-5\cos\zeta\cos\theta_{imp})$ (21)
$R_{cr}(\zeta,\theta_{imp})$ $=R_{cr,N}\left[1-\frac{9-\frac{18}{\pi}\theta_{imp}}{10}\frac{\cos\zeta}{2}\right]$ (22)
$k(\zeta,\theta_{imp})$ $=\frac{k_{N}}{2\pi}\exp(-0.02\cos\zeta\cos\theta_{imp}).$ (23)
In equation 22 the impact angle $\theta_{imp}$ is in radians.
The functions given by Raducan et al. (2022) in their equations 6–9 (and here with our equations 2023) are not consistent with their Figure 12 but appear consistent with the distributions shown in their other figures. The weak dependence of the function $k(\zeta,\theta_{imp})$ on angles $\theta_{imp}$ and $\zeta$ is supported by laboratory measurements of the azimuthal ejecta mass distribution from 104 m/s oblique impacts into sand (Quillen and Doran, 2024).
With the scaling functions for oblique impacts (Equations 20 - 22), we modify Equation 13, giving an oblique impact ejecta velocity function,
$\frac{v_{ej}(x,\pi_{4},\theta_{imp},\zeta)}{v_{pj}}=$ $\ C_{1}(\theta_{imp},\zeta)x^{-\frac{1}{\mu(\theta_{imp},\zeta)}}\pi_{4}^{-\frac{\nu}{\mu(\theta_{imp},\zeta)}}$
×$\ \times\left(1-\frac{x}{n_{2}R_{cr}(\theta_{imp},\zeta)}\right)^{p},$ (24)
where $\mu(\theta_{imp},\zeta),C_{1}(\theta_{imp},\zeta)$ and $R_{cr}(\theta_{imp},\zeta)$ are calculated from Equations 2022. Equation 15 for the mass distribution, is similarly modified to depend on azimuthal and impact angles
$\frac{M(<x,\pi_{4},\theta_{imp},\zeta)}{m_{pj}}=\frac{3k(\theta_{imp},\zeta)}{4\pi}\pi_{4}\left(\left(\frac{x}{a_{pj}}\right)^{3}-n_{1}^{3}\right)$ (25)
with $k(\theta_{imp},\zeta)$ from equation 23.
2.5 Ejecta Velocity Components
Figure 4: We illustrate how rotation affects the ejecta velocity as seen in an inertial frame. Black arrows show rotation, Brown shows ejecta particle velocity in the frame rotating with the surface at an angle from normal of about $45^{\circ}$. To transfer into the inertial frame, we add the rotational velocity, resulting in the purple arrows showing the ejecta velocity in the inertial frame. The purple arrows are longer on the trailing side, implying that angular momentum is carried away by ejecta. Ejecta curtains are computed in the frame rotating with the surface (brown arrows), however the likelihood of escape depends on whether the ejecta velocity in an inertial frame (purple arrows), exceeds the escape velocity.
Ejecta distributions are calculated in the frame rotating with the surface; however, to determine which ejecta escape the planetesimal, we compare the ejecta velocities in the inertial frame to the escape velocity. Because the maximum ejecta velocity is usually less than the projectile velocity, the ratio of ejecta velocity to rotational velocity is larger than the ratio of projectile velocity to rotational velocity. This means when we transfer back into the inertial frame, planetesimal rotation causes a more significant change in the ejecta velocity, than there was when we transferred into the rotating frame. This is illustrated in Figures 2 - 4. In other words, we could have neglected the rotational velocity when estimating the ejecta distribution, but should not neglect the rotational velocity when determining which ejecta escapes.
In the surface rotational frame, and in terms of the surface normal, downrange and binormal vector directions, the ejecta velocity is
${\bf v}_{ej}=v_{ej,d}\hat{\bf d}+v_{ej,b}\hat{\bf b}+v_{ej,n}\hat{\bf n}$ (26)
with components
$v_{ej,d}$ $=v_{ej}\cos(\zeta+\pi)\cos(\theta_{ej})$
$v_{ej,b}$ $=v_{ej}\sin(\zeta+\pi)\cos(\theta_{ej})$
$v_{ej,n}$ $=v_{ej}\sin(\theta_{ej}).$ (27)
Here $v_{ej,d}$ is the downrange component, $v_{ej,n}$ is the normal component, $v_{ej,b}$ is the binormal component, and $\theta_{ej}$ is the ejecta angle. Experiments and simulations find that the ejecta angle $\theta_{ej}$ varies by less than 10 degrees as a function of impact angle, launch position and azimuthal angle (Anderson et al., 2003; Raducan et al., 2022). Hence we assume that $\theta_{ej}$ is constant.
For each impact we compute ejecta distributions in the frame rotating with the surface using equations 24 for the velocity and 25 for the mass distribution. To determine which ejecta escape the planetesimal, we transfer the ejecta velocity back into the inertial frame by reversing equation 5;
${\bf u}_{ej}={\bf v}_{ej}+{\bf u}_{rot}.$ (28)
2.6 The Mass and Momentum of Escaping Ejecta
Following Holsapple and Housen (2012), we assume that a parcel of ejecta with velocity greater than the escape velocity, $u_{ej}\geq u_{esc}$, escapes from the system. We define a function that describes whether a parcel of ejecta escapes,
$V(x,\pi_{4},\theta_{imp},\zeta)=\begin{cases}1,&u_{ej}(x,\pi_{4},\theta_{imp},\zeta)\geq u_{esc}\\ 0,&u_{ej}(x,\pi_{4},\theta_{imp},\zeta)<u_{esc}\end{cases},$ (29)
where the escape velocity is from Equation 2.
To find the ejecta mass for a specific launch location $x$ from the site of impact, we differentiate Equation 15 (with respect to $x$) giving
$\frac{d\left(M(<x,\pi_{4},\theta_{imp},\zeta)\right)}{m_{pj}}=\frac{9k(\theta_{imp},\zeta)}{4\pi}\pi_{4}x^{2}dxd\zeta.$ (30)
Integrating this over possible launch position $x$ within the crater radius and for different azimuthal angles $\zeta$, we compute the total escaping ejecta mass from a single impact;
$M_{ej,esc}=\int_{x}\int_{\zeta}d\left(M(<x,\pi_{4},\theta_{imp},\zeta)\right)V(x,\pi_{4},\theta_{imp},\zeta).$ (31)
Since we neglect ricochet, we assume the mass of the projectile is either accreted onto the planetesimal or incorporated into ejecta. Thus $M_{ej,esc}/m_{pj}<1$ indicates accretion, while $M_{ej,esc}/m_{pj}>1$ indicates erosion.
For a distribution of impacts, we define a mass ratio
$\epsilon_{M}\equiv\biggl{<}\frac{M_{ej,esc}}{m_{pj}}\biggl{>}$ (32)
which is the average fraction of mass ejected per unit projectile mass.
We denote the momentum in escaping ejecta in the rotating frame as ${\bf p}_{ej,esc}$. We calculate the downrange component of the escaping ejecta
${\bf p}_{ej,esc}\cdot\hat{\bf d}=$ $\int_{x}\int_{\zeta}({\bf v}_{ej}\cdot\hat{\bf d})\ d\left(M(<x,\pi_{4},\theta_{imp},\zeta)\right)\times$
$\qquad\ \ V(x,\pi_{4},\theta_{imp},\zeta).$ (33)
The $\hat{\bf b}$ momentum component of the escaping ejecta can be neglected because the adopted ejecta distribution is symmetric about the $\hat{\bf d}$-$\hat{\bf n}$ plane. We don’t compute the normal momentum component because it does not contribute to the torque on the planetesimal.
2.7 The Angular Momentum Transfer Efficiency
The change in planetesimal angular momentum due to the impact at location ${\bf r}$ would be
$\Delta{\bf L}_{a}=m_{pj}{\bf r}\times{\bf u}_{pj}-{\bf r}\times{\bf p}_{ej,esc}$ (34)
repeating equation 67. See A for why this depends upon the momentum of escaping ejecta computed in the rotating frame and why we can neglect the ejecta trajectory after impact.
When averaging over a distribution of impacts that lack angular momentum with respect to the planetesimal, the projectile momentum term in equation 34 can be neglected as it averages to zero. We only need to calculate the rightmost term that depends on the momentum of escaping ejecta as seen in the rotating frame. Because the planetesimal is orbiting the star, it is likely to feel a headwind coming from different directions in the orbital plane (see Figure 1). We assume that $\Delta{\bf L}_{a}$ components that are not aligned with the planetesimal’s spin axis average to zero. This implies that the average change in angular momentum per impact for a distribution of projectiles is
$\langle\Delta{\bf L}_{a}\rangle=\left(\bigl{<}{\bf r}\times({\bf p}_{ej,esc}\cdot\hat{\bf{d}})\hat{\bf{d}}\bigl{>}\ \cdot\ \hat{\bm{\Omega}}_{a}\right)\hat{\bm{\Omega}}_{a}$ (35)
where $\hat{\bm{\Omega}}_{a}$ is the unit vector along the spin axis. We only need to use the downrange component of escaping ejecta (equation 33) to calculate the angular momentum parallel to the spin axis, as that is transferred to the planetesimal via the escaping ejecta. The change in angular momentum of the planetesimal per unit projectile mass is equal to $\langle\Delta{\bf L}_{a}\rangle/m_{pj}$.
We define a dimensionless angular momentum transfer efficiency
$\epsilon_{L}\equiv-\frac{\langle\Delta{\bf L}_{a}\rangle}{L_{a}}\frac{M_{a}}{m_{pj}}$ (36)
where $L_{a}=\frac{2}{5}M_{a}R_{a}^{2}\Omega_{a}$ is the planetesimal’s spin angular momentum and that of a homogeneous spinning sphere.
The angular momentum transfer efficiency is the fractional change in angular momentum caused by a distribution of projectiles per unit projectile mass in units of planetesimal mass. The amount of projectile mass required to remove a significant fraction of the planetesimal’s spin angular momentum would be approximately $M_{a}/\epsilon_{L}$.
Parameter Symbol Fiducial Value
Planetesimal Radius $R_{a}$ 10 000 [m]
Planetesimal Density $\rho_{a}$ 235 [kg m-3]
Planetesimal Strength Ya 500 [Pa]
Planetesimal Mass Ma $\frac{4\pi}{3}\rho_{a}R_{a}^{3}$ [kg]
Planetesimal Rotation Velocity $\Omega_{a}$ $\Omega_{\rm break-up}$
Headwind Velocity (Inertial) $u_{hw}$ 30 [m/s]
Projectile Density $\rho_{pj}$ 235 [kg m-3]
Ejecta Angle $\theta_{ej}$ 45
Table 1: Planetesimal fiducial parameters assumed throughout the simulations unless other values are stated. These parameters are based on the Kuiper Belt Object (486958) Arrokoth (Keane et al., 2022; Brisset et al., 2022), with the radius being the spherical equivalent from the mass and density of Arrokoth. The projectile velocity is taken to be a value near the headwind velocity experienced by planetesimals in the outer disk (Quillen et al., 2024). For these fiducial parameters, the escape velocity from the planetesimal is 3.6 m/s.
Crater Scaling Parameters
$C_{1,N}$ $k_{N}$ $\nu$ $n_{1}$ $n_{2}$ $H_{2}$ $p$ $q$
0.55 0.3 0.4 1.2 1.0 0.4 0.3 0.2
Table 2: Crater scaling coefficients are taken from Table 3 by Housen and Holsapple (2011). We adopt parameters typical of granular materials.
2.8 Gravitational Focusing
For higher mass planetesimals, those with escape velocity above the projectile velocity, the gravitational field of the planetesimal causes gravitational focusing. In this case the velocity and angle of impact is sensitive to the orbit of the projectile when approaching the planetesimal. We compute the velocity and location of the impact using a hyperbolic orbit, neglecting gas drag. To collide with a spherical object of mass $M_{a}$ and radius $R_{a}$, a projectile must have impact parameter $b$ lower than
$b_{max}=R_{a}\sqrt{1+\frac{2GM_{a}}{u_{hw}^{2}R_{a}}}.$ (37)
We have neglected the size of the projectile, assuming that it is small compared to the planetesimal. Here $u_{hw}$ is the headwind velocity distant from the planetesimal. We assume that the projectile has the same velocity. With an impact parameter of $b_{max}$, the projectile would graze the planetesimal.
We define a new Cartesian coordinate system with coordinates $(x_{o},y_{o},z_{o})$ to describe the hyperbolic orbit of the projectile. The orbit of the projectile lies in the $x_{o},y_{o}$ plane. The angle $\alpha$ is the angle between the projectile orbital plane and the xy plane of the circumstellar disk. We assume that the projectile arrives from the $x_{o}=+\infty$ direction. We determine the components of the impact location on the planetesimal surface
${\bf r}_{impact}=R_{a}(\cos(\Delta f)\hat{\bf x}_{o}+\sin(\Delta f)\hat{\bf y}_{o}).$ (38)
where $\Delta f$ is the difference between the maximum true anomaly and the true anomaly at impact. See B, and specifically equation 73, for an expression for $\Delta f$.
We calculate the impact velocity in the inertial frame using energy conservation,
$|\textbf{u}_{pj}|=\sqrt{u_{hw}^{2}+\frac{2GM_{a}}{R_{a}}}.$ (39)
Only when we neglect gravitational focusing entirely is the velocity at impact equal to that of the headwind, $u_{pj}=u_{hw}$. Even on low mass planetesimals, with $u_{hw}<u_{esc}$, the gravitational field would affect the impact velocity $u_{pj}$. The collision angle (the angle between the surface normal and the orbital trajectory at the impact point) is
$\gamma_{c}={\rm atan}\left(\frac{1+e\cos f_{impact}}{e\sin f_{impact}}\right)$ (40)
with the convention that a grazing impact has $\gamma_{c}=\frac{\pi}{2}$. See B Equation 71 for an equation for the true anomaly of the impact, $f_{impact}$. The impact velocity vector in the inertial reference frame is then
${\bf u}_{pj}=-|u_{pj}|\left(\cos(\Delta f-\gamma_{c})\hat{\bf x}_{o}+\sin(\Delta f-\gamma_{c})\hat{\bf y}_{o}\right).$ (41)
We calculate the projectile velocity using Equation 41 for all impacts, regardless of the planetesimal mass.
To go into the $x,y,z$ coordinate system relevant for computing the ejecta distributions, shown in Figure 1, (taking into account the orbit orientation) we rotate the calculated location of impact and the velocity of impact around the x-axis with rotation matrix
$R(\alpha)_{x}=\begin{pmatrix}1&0&0\\ 0&\cos\alpha&-\sin\alpha\\ 0&\sin\alpha&\cos\alpha\\ \end{pmatrix}.$ (42)
The resulting impact velocity is used to compute the impact angle in the rotating frame, with Equation 9.
2.9 Caveats
While computing the integrated erosion rate and the angular momentum transfer efficiency in this section, we have neglected some degrees of freedom. We mention these points here as possible directions of future study.
2.9.1 Gravitational Focusing
We have neglected gravitational focusing of the escaping ejecta particles. Neglecting gravitational ”defocusing” does not influence our estimate of the angular momentum transfer efficiency, but would influence computation of the linear momentum transfer efficiency.
2.9.2 Planetesimal Shape and Structure
For simplicity, we have neglected the shape of the planetesimal by assuming that it is spherical. If we allow for non-spherical planetesimals, the gravity vector at a surface location is not necessarily perpendicular to the surface, so the surface can have a slope with respect to the direction of self-gravity. Studies measuring crater volume, diameter, and axis ratio of oblique impacts into inclined granular media find that ejecta distributions also depend on surface slope (Takizawa and Katsuragi, 2020; Omura et al., 2021). In future, models could be extended to take into account impact angle, surface slope and the direction and values of ejecta velocity vectors that allow ejecta to escape. Such models could find an analogies to the Yarkosky effect, causing orbital drift, but mediated by momentum exchange with a distribution of projectiles.
We neglected heterogeneity of the planetesimal surface. Its surface could be rough and contain particles of different sizes, densities, compositions, and shapes. These variations are not accounted for in the ejecta distribution models that are used throughout this paper but might influence the fraction and momentum content of escaping ejecta.
2.9.3 Gas Drag Around Planetesimals
Crater ejecta would be effected by the headwind flowing around the surface of the planetesimal, as shown by Quillen et al. (2024). Accurate models of rarefied gas flow around a planetesimal would help determine the quantity of ejecta that escapes, as small ejecta particles are swept up by a disk wind near the planetesimal surface. A wind would accelerate small particles in low velocity ejecta to beyond the escape velocity. However, the fraction of material swept up would be approximately the same equal on leading and trailing sides of the planetesimal, so we suspect that the total contribution to the angular momentum loss would be low.
2.9.4 Rotating Projectiles
In this study we have neglected projectile spin. However, studies of impacts from spinning projectiles (Carvalho et al., 2023a, b, 2024) find that crater profiles are sensitive to projectile spin. To accurately take into account the projectile spin, we would need an accurate description of the spin angular momentum distribution of the particles within the source of projectiles. Additionally, crater scaling coefficients for the ejecta could become dependent upon the projectile spin.
3 Ejecta Distributions for Single Impacts
Figure 5: We show a polar plot of an impact with projectile density $\rho_{pj}=1000$ kg m-3 occurring at a Latitude = 0, Longitude = 0. We define Longitude = 0 as the furthest point on the headwind-facing hemisphere. (a) In the frame moving with the surface, we show ejecta velocity in color as a function of ejecta launch position and azimuthal angle. The radius on the plot refers to distance from impact point (in units of projectile radius) and the polar angle on the plot is $\zeta$, the azimuthal angle. Due to the rotation of the planetesimal, the impact angle in the frame rotating with the surface of this collision is $\theta_{I}\approx 87.1^{\circ}$ despite being a head-on collision. (b) Similar to Figure 5a, except we only show pixels that are above the escape velocity, highlighting regions of the crater that lead to angular momentum transfer and mass loss. Despite being a head-on collision, planetesimal rotation causes an asymmetric ejecta curtain.
Figure 6: We show a polar plot of the ejecta distributions as a function of ejection position for an impact with impact parameter $b=0.75R_{a}$ and orbit angle $\alpha$ = 0, on a 20 km-radius planetesimal with and without gravitational focusing. The radius on these plots is distance from impact point and the angle, the azimuthal angle $\zeta$. (a) We show the fractions of mass ejected above escape velocity. In this impact, gravitational focusing is neglected and the impact velocity is equal to that at large distances from the planetesimal $u_{pj}=u_{hw}=30$ m/s. The impact occurs at surface coordinates (lat $=0^{\circ}$, long $\sim 48.6^{\circ}$), with impact angle $\theta_{imp}\sim 48.8^{\circ}$. (b) Similar to a) except gravitational focusing is taken into account. The impact occurs at (lat $=0^{\circ}$, long $\sim 47.5^{\circ}$) with $\theta_{imp}\sim 50.6^{\circ}$ with projectile velocity $u_{pj}\sim 30.86$ m s-1. (c) Residuals of ejecta mass for the difference of (b) - (a) in units of projectile mass. (d) We show the difference in ejecta velocity in units of the escape velocity, $u_{esc}$, for the impact with gravitational focusing minus that omitting gravitational focusing. Despite the fact that the escape velocity is lower than the projectile speed distant from the planetesimal, the residuals are not small. Residuals are largest in the downrange direction.
In Section 2 we explained how we estimate the total ejecta mass that escapes and the angular momentum transferred to the planetesimal from a distribution of impacts. In this section we examine the ejecta distributions from a single impact. Additionally, we will explore the role of gravitational focusing.
Unless otherwise stated, we assume a set of fiducial values for the planetesimal and projectile properties, which are listed in Table 1. Particle ejecta angles (measured from the surface normal) tend to lie between 30 - 60 with a narrow variation (10 - 20) (Stoeffler et al., 1975; Anderson et al., 2003, 2004; Tsujido et al., 2015; Wünnemann et al., 2016; Luther et al., 2018; Raducan et al., 2022; Suo et al., 2024). We adopt a constant 45 ejecta angle.
For each impact, we generate a 2 dimensional polar grid of ejecta launch positions ($x$ values) and azimuthal angles ($\zeta$ values). Here $x$ is the radial position within the crater centered at the site of impact impact (see Figure 3), in units of the projectile radius, $a_{pj}$. At each point on this grid we compute an ejecta velocity velocity in the rotating frame, an ejecta velocity vector in the inertial frame, and the ejecta mass. We calculate the crater radius $R_{cr,N}$ for a normal impact using Equation 14. Following the point-source approximation, we reject locations with $x<n_{1}a_{pj}$ or $x\leq n_{2,N}R_{cr}$ by setting the ejecta mass and velocity at those grid points to zero. For all remaining $x$ and $\zeta$ points within our grid, we calculate the ejecta velocity and mass. The ejecta velocity $v_{ej}$ in the rotating frame within an oblique impact crater is given by Equation 24, and crater radius $R_{cr}$ in equation 22.
Using Equations 26 and 27, we calculate the ejecta velocity components in the rotating frame at each point on our grid. We also compute the inertial frame ejecta velocity ${\bf u}_{ej}$ (via Equations 26). The ejecta velocity is compared to the escape velocity. We assume that particles escape if $u_{ej}\geq u_{esc}$.
Figure 5 shows an impact that, absent rotation, would be a normal impact at longitude (long) = 0, latitude (lat) = 0, where we define long = 0 to be the furthest point of the planetesimal on the headwind-facing hemisphere. The influence of rotation can be seen by the fact that the ejecta curtain is not azimuthally symmetric, but rather it is oblong because it is an oblique impact, with impact angle $\theta_{imp}\approx 86.61^{\circ}$. The figure show a polar plot with angle equal to the azimuthal angle $\zeta$ and distance between the origin to the ejecta launch distance $x$ from the impact site in units of projectile radius, $a_{pj}$. The color shows the ejecta velocity (in the rotating frame) which decreases as a function of ejecta launch distance.
Because the impact in Figure 5 is a nearly normal impact, the projectile velocity is nearly but not exactly azimuthally symmetric, as shown in Figure 5a. In Figure 5b, the color also shows ejecta velocity in the rotating frame, but we only show points with ejecta velocity in the inertial frame that lie above the escape velocity. The boundary is not azimuthally symmetric. We find that the rotation is important in determining which ejecta escapes, even at low rotation velocities.
3.1 Importance of Gravitational Focusing
In this section we compare ejecta distributions from a single impact assuming no gravitational deflection with one that is deflected on a hyperbolic orbit. The projectile has density $\rho_{pj}=1000$ kg m-3, has impact parameter $b=0.75R_{a}$ and $\alpha=0$ (describing the orbital tilt). The distant headwind velocity is equal to the fiducial value of $u_{hw}=30$ m s-1. The planetesimal radius in this example is $R_{a}=20$ km and barely large enough that the orbital deflection could be important. Using $\rho_{a}=235$ kg m-3 and Equation 2, we find that the escape velocity from the planetesimal $u_{esc}\approx 7.24$ m s-1.
Figure 6 compares the two ejecta distributions. Figure 6 a and b show the ejecta mass distributions, with and without gravitational focusing, as a function of ejection site in polar coordinates. Consequently, the locations and velocities of the two impacts differ. In Figure 6c, we subtract the ejecta mass distributions from each impact and show the mass residuals in units of $d(M(x,\pi_{4},\theta_{imp},\zeta)/m_{pj})$. As a percentage, the residual shows a maximum change of $\sim$4in the ejecta mass distribution as a result of gravitational focusing. In Figure 6d, the velocity distributions are subtracted and the residual is plotted in units of escape velocity, $u_{esc}$. The residuals for ejecta velocity are large, particularly for the highest velocity material that is likely to escape the planetesimal, with the greatest residual equating to a $\sim$12change.
The inclusion of gravitational focusing results in a change to all of the impact parameters calculated. The impact velocity for the non-deflected case is equal to the distant headwind velocity, $u_{pj}=u_{hw}$ (30 m s-1), whereas in the deflected case $u_{pj}\sim$ 30.86 m s-1. The impact locations differ slightly between the non-deflected and deflected impacts, (lat = 0, long $\sim 48.6^{\circ}$) and (lat $=0^{\circ}$, long $\sim 47.5^{\circ}$), respectively. Similarly, the two impacts have different impact angles for the non-deflected and deflected impacts, ($\sim 48.8^{\circ}$ vs. $\sim 50.6^{\circ}$). The differences in impact angle (a result of different impact locations) and projectile velocity caused by gravitational focusing results in the changes in the ejecta distributions seen in Figure 6.
For a planetesimal with radius of $R_{a}=20$ km we see the influences of gravitational focusing altering the ejecta distributions due to changes in the impact parameters. For a planetesimal with this radius (20 km), the maximum impact parameter (Equation 37) is $b_{max}\sim 20575$ m. Thus, for a projectile with a $b/b_{max}\approx$ 0.729 (Due to assuming b = 0.75 $R_{a}$ as was done in Figure 6), and a planetesimal with $R_{a}/b_{max}\approx$ 0.97, the effects of gravitational focusing are no longer negligible. Mild deflection of a projectile trajectory due to a low-mass planetesimal results in a noticeable change to the impact angle, velocity, and location, resulting in changes of a few percent to the ejecta distribution for the projectile.
3.2 Caveat: Ejecta Models
In our integrations we use an ejecta angle of 45; however, a distribution of angles sensitive to azimuthal angle, impact angle, and launch position would allow for more physically accurate integrations. In future, functions for ejecta scaling parameters could be fit to simulated ejecta angle distributions at a range of impact velocities (Raducan et al. 2022 fit 4 functions but not the ejecta angle and only simulated high velocity impacts) or to experimental data (Anderson et al. 2003 measured this angle experimentally in oblique impacts but did not fit functions to them and also only considered high velocity impacts). While Quillen and Doran (2024) studied oblique impacts at intermediate impact velocity, they only studied the sensitivity of the integrated ejecta mass and crater radius to impact and azimuthal angle. There is a need for more experimental characterization of oblique impacts.
Cohesion is incorporated into a strength parameter describing the asteroid surface. However, this description may not capture granular dynamics, particularly if projectiles are similar in size to particles on the planetesimal surface.
Porosity can affect crater shape and ejecta properties (Housen and Holsapple, 2003), which we did not take into account. Variation due to porosity increasingly affects large craters when looking at small bodies, resulting in a reduction in ejecta mass, while for small craters the ejecta blankets are unaltered. When the product of gravitational acceleration and crater diameter, $g_{a}D_{c},\sim 10^{5}$ (in cgs), for planetesimal porosity of 42% and 70%, the ejected crater mass begins to level out and eject 80-90% the mass of an equivalent non-porous crater (Housen and Holsapple 2003).
4 Torque from a Distribution of Impacts
In the previous section we integrated the ejecta mass and velocity distributions for a single impact. In this section we integrate a distribution of impacts to estimate the total mass ejected and the angular momentum transferred to the planetesimal by a population of impactors.
To choose impact locations on the planetesimal surface, we generate a uniform distribution of coordinate pairs, $(x_{d},y_{d})$, and reject point pairs with an impact parameter $b>b_{max}$, noting that $b=\sqrt{x_{d}^{2}+y_{d}^{2}}$. The orbit angle $\alpha=\arctan(y_{d}/x_{d})$, and impact parameter $b$ determine the impact location, velocity and angle, as outlined in B and the maximum impact parameter $b_{max}$ is given in equation 37.
To model impacts from particles moving in a headwind, we assume that far from the planetesimal the projectiles are all unidirectional with magnitude equal to the headwind velocity, $u_{hw}$. We take gravitational focusing into account for all planetesimal masses. Consequently we use hyperbolic orbits to calculate the projectile impact location, velocity, and impact angle on the planetesimal surface, as described in Section 2.8.
The collisions are integrated to find the mass loss ratio (see Equation 36) and angular momentum transfer efficiency (as defined in section 2.7, Equations 31, 33 - 36). We perform similar integrations but with different planetesimal and projectile properties to understand how these properties affect the angular momentum transfer efficiency and mass loss ratio.
Figure 7: Mass loss ratio and angular momentum transfer efficiency for a planetesimal with the fiducial values from Table 1 as a function of planetesimal radius on the x-axis. In each panel pair (a-b, c-d, e-f, g-h), the effect of varying a single parameter is shown. Panels (a,b) varies $\pi_{4}$, (b,c) varies planetesimal strength $Y_{a}$, (e,f) varies the distant headwind velocity $u_{hw}$, and (g,h) varies the planetesimal spin rate. The left column (subfigures a,c,e,g) show the angular momentum transfer efficiency and the right column (subfigures b, d, f, h) show the mass loss ratio. In the mass loss ratio plots b, d, f, h, we indicate regions of net accretion in green shading and net erosion in red shading. We find that the greatest influence on the angular momentum transfer are due to $\pi_{4}$ (the ratio of planetesimal to projectile density) and headwind velocity $u_{hw}$, while the mass loss ratio is most affected by the planetesimal strength $Y_{a}$ and head wind velocity $u_{hw}$.
In Figure 7, we plot the angular momentum transfer efficiency, $\epsilon_{L}$ and the mass loss ratio $\epsilon_{M}$ as a function of planetesimal radius $R_{a}$. All quantities, when not specified, are taken to be those of the fiducial parameters listed in Table 1. In Figure 7a and 7b, we vary the projectile density $\rho_{pj}$ giving different $\pi_{4}$ values. These panels show how the density ratio between the projectile and planetesimal affect the mass loss ratio and angular momentum change as a function of planetesimal radius $R_{a}$. Figures 7c - 7h show how variations in strength $Y_{a}$, headwind velocity $u_{hw}$, and spin $\Omega_{a}$ affect the mass loss ratio and angular momentum transfer efficiency.
The panels in the left column of Figure 7 show that the angular momentum transfer efficiency is positive, confirming that the mechanism proposed by Dobrovolskis and Burns (1984) operates in a low velocity regime. However, the angular transfer efficiency is low, at most $\epsilon_{L}\sim$ 0.07 in the most extreme case at $\pi_{4}=0.25$. We find that the mechanism for spin down is inefficient. The panels in the right column show that, for our fiducial parameters, the total ejected mass remains in an accretionary regime, with the accretion being more efficient for larger radii planetesimals which have higher escape velocities. Thus the angular momentum drain mechanism proposed by Dobrovolskis and Burns (1984) also operates in an accretionary impact regime.
In Figure 7, for planetesimal radius $R_{a}\lesssim 10$ km, the angular momentum transfer efficiency $\epsilon_{L}$ is nearly independent of planetesimal radius. Above $R_{a}\sim 10$ km the angular momentum transfer efficiency tends to decrease, and this we attribute to due to gravitational focusing increasing the projectile impact velocity.
In Figure 7c - 7d we see low-strength planetesimal materials exhibit marginally increased angular momentum transfer efficiency at large planetesimal radii. At lower mass and radii, the effect flips, with low strength planetesimals exhibiting the least angular momentum spindown, with an intermediate strength values being most effective in this regime. At all points, especially at lower mass and radii, lower strength bodies exhibit increased mass loss because each crater sheds more material from the planetesimal when compared to one on a higher strength planetesimal.
In Figure 7e-7f we show how the headwind velocity affects the mass loss ratio and angular momentum transfer efficiency. High wind velocities result in an increase in mass loss ratio, $\epsilon_{M}$, and a decrease in angular momentum transfer efficiency, $\epsilon_{L}$. At low headwind velocities, the ejecta velocity is smaller, thus planetesimal rotation has a greater influence in determining which ejecta escapes, as illustrated by Figure 4. This leads to an increase in escaping ejecta on the trailing hemisphere, leading to larger spin-down. At higher headwind velocities, the ejecta velocity is greater and we find more ejecta escaping downrange on the leading hemisphere, leading to a lower overall spin-down.
In Figure 7g-7h, we show sensitivity to the planetesimal spin. We find that the accretion rate is insensitive to the spin rate, and the angular momentum transfer efficiency decreases as a function of decreasing spin rate.
For all but one variable we see the changes to the mass loss ratio and angular momentum transfer efficiency are aligned, we see an increase/decrease in both effects across all planetesimal masses. However, the headwind velocity exhibits the opposite trend, with an increase in $\epsilon_{L}$ associated with a decrease in $\epsilon_{M}$. To explore why this is true, we create a new figure to investigate the asymmetric influence of planetesimal rotation on the angular momentum transfer efficiency for the leading vs trailing hemispheres.
In Figure 8, we show the angular momentum transfer efficiency in longitudinal bins across the planetesimal, denoted as $\epsilon_{L,\phi}$. The summation of all $\epsilon_{L,\phi}$ results in $\epsilon_{L}$, which we have been plotting in Figure 7. In each panel we show histograms at two different headwind velocities, and the difference in the two histograms is shown as a red line which we label ‘residual’. The darker histogram shows the higher headwind velocity. When the headwind velocity is higher, the two sides of the histogram are more symmetrical and this reduces the angular momentum lost in ejecta, even though more ejecta escapes from higher velocity impacts. This figure is consistent with the reduction in angular momentum transfer efficiency at high headwind velocities seen in Figure 7e,f. In Figure 9, is similar to Figure 8 except we show longitudinal bins for two integrations with different $\pi_{4}$ values. The asymmetry evident in Figure 9 is absent, indicating the increase in $\epsilon_{L}$ associated with a decrease in $\epsilon_{M}$ while varying $u_{hw}$.
Among the varied parameters $\pi_{4},Y_{a},u_{hw},\Omega_{a}$ Figure 7a shows that the angular momentum transfer is most sensitive to the ratio of planetesimal to projectile density, $\pi_{4}$. A high density projectile colliding with a low density planetesimal will yield significantly more ejecta and so more angular momentum transfer than when the densities are equal.
Figure 8: We plot a histogram of the angular momentum transfer efficiency contribution from individual longitudinal bins across the planetesimal surface. In each panel, we plot histograms at two headwind velocities, $u_{hw}$. The left side of the histograms are taller and this is why spin angular momentum is carried by the ejecta. In each panel we show histograms at two different headwind velocities and the difference in the two histograms is shown as a red line which we label ’residual’. The darker histogram shows the higher headwind velocity. When the headwind velocity is higher, the two sides of the histogram are more symmetrical and this reduces the angular momentum lost in ejecta, even though more ejecta escapes from higher velocity impacts. This figure is consistent with the reduction in angular momentum transfer efficiency at high headwind velocities seen in Figure 7e,f. (We indicate that the Trailing Hemisphere, as shown in Figure 2, lies between $-100^{\circ}<\phi<0^{\circ}$, and the Leading Hemisphere lies between $0^{\circ}<\phi<100^{\circ}$)
Figure 9: We plot histograms of the contribution to the angular momentum transfer efficiency from longitudinal bins across the planetesimal surface, similar to Figure 8, except here histograms with different density ratios $\pi_{4}$ are shown. When $\pi_{4}$ is varied, both left and right sides of the histogram vary in height, whereas in Figure 8, only one side was affected.
4.1 A Simple Velocity Gradient
A planetesimal embedded in a disk that is undergoing streaming instability is not an isolated object traversing through a constant headwind. A pebble cloud from which a planetesimal forms would have some inherent angular momentum of its own (Nesvorný et al., 2021; Lorek and Johansen, 2024). The streaming instability generates extended concentrations of dust particles known as filaments (Johansen et al., 2007; Schäfer et al., 2017; Johansen et al., 2015). We consider a gravitationally bound pebble cloud that is within a filament which has a velocity gradient giving an angular momentum component to the projectile velocity distribution.
We use the angular momentum of a gravitationally bound pebble cloud to estimate an upper bound for a spatial velocity gradient. If we assume that the pebble cloud extends to the hill radius, $R_{cloud}=R_{H}$, where $R_{H}=a\big{(}M_{cloud}/(3M_{*})\big{)}^{\frac{1}{3}}$. We take the material within the pebble cloud to be orbiting at breakup velocity, using Equation 3 replacing the planetesimal density, $\rho_{a}$, with the modified hill density, $\rho_{H*}$. We find that the pebble cloud angular momentum is
$L_{cloud}=\frac{2}{5}M_{cloud}R_{cloud}^{2}\Omega_{cloud}$
$\sim 1.8\times 10^{-5}M_{cloud}^{5/3}\ {\rm kg^{-2/3}\ m^{2}\ s^{-1}}\left(\frac{a}{1{\rm AU}}\right)^{1/2}\left(\frac{M_{*}}{M_{\odot}}\right)^{-1/6}.$ (43)
We estimate that the filament’s angular momentum is related to a deviation in the headwind velocity across the filament $\Delta u_{hw}$.To determine the velocity gradient across the filament, we use the angular momentum density within the filament fluid flow with the dust volume density due to the increased particle density within the streaming instability (Li and Youdin, 2021). As such, the angular momentum density within the filament is
$l=\rho_{H*}({\bf r}_{f}\times\Delta u_{hw}).$ (44)
Approximating the filament as a cylinder, we find the angular momentum within the filament to be
[…]
Assuming the circular cross-section of the cylinder is aligned with the yz-plane, with the length of the cylinder and the direction parallel to the headwind velocity being aligned with $\hat{x}$. We assume a linear velocity gradient, $\Delta u_{hw}=\gamma y\hat{x}$, and denote ${\bf r}_{f}=y\hat{y}+z\hat{z}$. We use $\gamma$ to determine the magnitude of the linear velocity gradient. We take the length of cylinder, $h$, required to contain a total mass of $M_{cloud}$.
${\bf L}_{filament}\sim 17.6\ \gamma M_{cloud}^{5/3}\left(\frac{a}{1{\rm\ AU}}\right)^{2}\left(\frac{M_{*}}{M_{\odot}}\right)^{-2/3}\ \hat{z}.$ (46)
Using Equations 43 & 46, we solve for the magnitude of the linear velocity gradient, $\gamma$:
$\gamma\sim 1\times 10^{-6}\ {\rm s^{-1}}\left(\frac{a}{1{\rm AU}}\right)^{-3/2}\left(\frac{M_{*}}{M_{\odot}}\right)^{1/2}.$ (47)
At an orbital distance of 45 AU, we approximate a velocity gradient of $\gamma\sim 3.3\times 10^{-9}\ {\rm s^{-1}}$ for a solar-mass star. The maximum change in headwind velocity, $|\Delta u_{hw}|$, experienced by a planetesimal will be at the maximum impact parameter, $b_{max}$,
$|\Delta u_{max}|$ $=\gamma b_{max}\sim 3\times 10^{-5}\ {\rm m\ s^{-1}}\times$
$\left(\frac{M_{cloud}}{M_{a}}\right)\left(\frac{a}{45{\rm AU}}\right)^{-3/2}\left(\frac{M_{*}}{M_{\odot}}\right)^{-13/6}\left(\frac{R_{a}}{10\ {\rm km}}\right)\times$
$\sqrt{1+\frac{8\pi G}{3}\left(\frac{\rho_{a}}{235\ {\rm kg\ m^{-3}}}\right)\left(\frac{u_{hw}}{30\ {\rm m\ s^{-1}}}\right)^{-2}\left(\frac{R_{a}}{10\ {\rm km}}\right)^{2}}$ (48)
which can be described as a fraction of the headwind velocity, $u_{hw}$,
$\frac{|\Delta u_{max}|}{u_{hw}}\sim 1\times 10^{-6}.$ (49)
The inclusion of a velocity gradient, with lower projectile velocities on the leading hemisphere and greater velocities on the trailing hemisphere, necessitates that the angular momentum of a distribution of impacts no longer has a mean of zero. Now, the projectile distribution contains angular momentum with respect to the planetesimal, and the angular momentum transfer from the projectiles no longer averages to zero. We include the first term of Equation 34, modifying the average change in angular momentum per impact for a distribution of particles;
$\langle\Delta{\bf L}_{a}\rangle=m_{pj}$ $\left(\bigl{<}{\bf r}\times{\bf u}_{pj}\bigl{>}\ \cdot\ \hat{\bm{\Omega}}_{a}\right)\hat{\bm{\Omega}}_{a}+$
$\left(\bigl{<}{\bf r}\times({\bf p}_{ej,esc}\cdot\hat{\bf{d}})\hat{\bf{d}}\bigl{>}\ \cdot\ \hat{\bm{\Omega}}_{a}\right)\hat{\bm{\Omega}}_{a}.$ (50)
The change in angular momentum of the planetesimal per unit projectile mass is still equal to $\langle\Delta{\bf L}_{a}\rangle/m_{pj}$.
We show the maximum change in angular momentum transfer efficiency and mass loss ratio due to the inclusion of an upper bound spatial velocity gradient in Figure 10. Despite looking at the upper bound for a velocity gradient within the projectile population, no detectable changes are calculated for the angular momentum transfer efficiency, $\epsilon_{L}$, and the mass loss ratio, $\epsilon_{M}$.
Simulations of collapsing pebble clouds indicate that the angular momentum of the pebble cloud range between 0.5 - 0.01 in hill units ($L_{H}=MR_{H}^{2}\sqrt{GM_{\odot}/a^{3}}$)(Nesvorný et al., 2021). Converting Equation 43 to be in units of the hill angular momentum we find $L_{cloud}\sim 2M_{cloud}R_{H}^{2}\sqrt{GM_{\odot}/a^{3}}$. Thus, assuming the upper bound of a pebble cloud rotating at breakup velocity results in a value between 4×200×4$\times$-200$\times$4 × - 200 × greater than the values seen in simulations. Due to there being no meaningful change to the angular momentum transfer efficiency and mass loss ratio at our assumed upper bound, we find it unnecessary to take into account the velocity gradient in a projectile population that would be required to form the pebble cloud.
Figure 10: Angular Momentum Transfer Efficiency and Mass Loss Ratio for a planetesimal and projectile with fiducial values shown in Table 1. The dotted line shows the values when including a spatial velocity gradient as determined by Equation 47. The solid line indicates the values when no velocity gradient is considered. Assuming the upper boundary for a velocity gradient remains incapable of meaningfully altering the angular momentum transfer efficiency and mass loss ratio.
4.2 Effective Spindown Over the Lifetime of the Collisional Era
The spindown rate of a planetesimal is dependent on the amount of projectile mass colliding with the planetesimal, thus we need an estimate of the total mass in disk or wind particles that collide with a planetesimal in a given time frame. We characterize the fraction of the protostellar disk mass in particles with the ratio $f_{p}\equiv\rho_{p}/\rho_{g}$, where $\rho_{p}$ is the particle density in the disk (the mass per unit volume in particles), not the density of the particles themselves, and $\rho_{g}$ is the gas density. We use a dimensionless factor $\xi_{p}$ to describe the mass fraction of disk particles that can impact a planetesimal. Taking into account only gravitational focusing, $\xi_{p}=b_{max}^{2}/R_{a}^{2}$. This factor could be lower as as some particles could be swept away by gas drag.
We adopt the circumstellar disk model from Quillen et al. (2024) which has the power law form for temperature and surface density of the minimum mass solar nebula by Hayashi (1981)
$\Sigma_{g}$ $=g_{\Sigma}\left(\frac{r}{10\text{AU}}\right)^{-3/2}$ (51)
$T_{g}$ $=g_{T}\left(\frac{r}{1\text{AU}}\right)^{-1/2}$ (52)
where $g_{\Sigma}$ and $g_{T}$ are coefficients that can be used to adjust disk surface density and temperature. For this disk model the headwind velocity is independent of radius and $u_{hw}\approx 38$ m/s.
Equation by Quillen et al. (2024) estimates the mass of particles from a disk impacting the planetesimal per unit area and time at the headwind speed
$F_{m}$ $\sim f_{p}\xi_{p}\rho_{g}u_{hw}$
$\approx 8\times 10^{-9}\ {\rm kg~{}m}^{-2}{\rm s}^{-1}\left(\frac{f_{p}}{10^{-2}}\right)$
×$\ \ \ \times\left(\frac{r}{10\ {\rm AU}}\right)^{-\frac{11}{4}}\frac{b_{max}^{2}}{R_{a}^{2}}g_{\Sigma}g_{T}^{\frac{1}{2}}.$ (53)
Assuming a planetesimal is embedded in disk undergoing streaming instability for $10^{3}$ years at 10 AU, with $f_{p}\sim 10^{2}$ (as streaming instability concentrates the particles), we estimate a total mass in projectiles
$\frac{\Delta M_{a}}{M_{a}}$ $=\frac{4\pi R_{a}^{2}F_{m}\Delta t}{M_{a}}$ (54)
$\sim 3\ \left(\frac{f_{p}}{10^{2}}\right)\left(\frac{r}{10\ {\rm AU}}\right)^{-\frac{11}{4}}g_{\Sigma}g_{T}^{\frac{1}{2}}\left(\frac{\rho_{a}}{235\ {\rm kg~{}m}^{-3}}\right)^{-1}$
×$\ \ \ \ \times\left(\frac{\Delta t}{10^{3}\ {\rm yr}}\right)\left(\frac{R_{a}}{10\rm km}\right)^{-1}\left(1+\frac{2GM_{a}}{u_{hw}^{2}R_{a}}\right).$ (55)
During an epoch of streaming instability, a recently formed planetesimal could be impacted by about its own weight in pebbles. Nevertheless because escaping ejecta is inefficient at removing spin angular momentum, at most a few percent of its spin could be carried away by ejecta even with this high impact rate.
In this section we found that the total mass in impacts experienced by an embedded planetesimal could be large compared to its total, particularly if the planetesimal is small. In our study we have found that impacts can drain a planetesimal’s angular momentum, however an impact more conventionally deflects an object. Computation of the momentum transfer efficiency would require taking into account the different directions that ejecta escape the planetesimal, during gravitational defocusing, as discussed in A. As the ejecta velocity is slower than the impact velocity, gravitational defocusing could be an even stronger effect on the escaping ejecta than on the incoming projectile.
The momentum change caused by a distribution of impacts is likely in the downwind direction, akin to a drag force. If the planetesimal is part of binary, impact induced drag would probably aid in removing angular momentum from the binary, causing it to tighten or merge. The large pebble mass flux present during streaming instability might facilitate binary merging. Hence it might be interesting in a future study to leverage ejecta scaling to compute the momentum transfer efficiency.
4.3 Caveat: Impactor Density Distribution
Throughout our integrations, we assumed all collisions occurred with a single projectile density $\rho_{pj}$. Since the ratio of projectile to planetesimal density, $\pi_{4}$, affects the integrated erosion and angular momentum transfer efficiency (See Section 4), the impactor density distribution would affect the total spin-down observed, due to the total projectile mass not all being of the same density. As such, evolutionary models on the size and density distribution of disk particulates (e.g. Colmenares et al. 2024) could be used to expand our extend our models.
5 Discussion
We discuss our work in the context of the study by Nesvorný et al. (2021), who studied KBO binary formation with N-body soft sphere (PKDGRAV). Nesvorný et al. (2021) found that the angular momentum of a pebble cloud prevents the rapid collapse into a single planetesimal. Instead, binary systems were more likely to form, with only a fraction of the pebble cloud being incorporated into the planetesimals. In this section we consider the fraction of pebble cloud angular momentum that must be removed to collapse a single planetesimal and place this value in context with the fraction of angular momentum that can be removed by escaping ejecta.
5.1 The Angular Momentum of Planetesimals Formed Via Streaming Instability
Nesvorný et al. (2021) investigated 3 streaming instability simulations in which clumps formed. They measured the angular momentum of these clumps and computed a ratio they called the mass scaled angular momentum $l=L_{c}/L_{H}$ where $L_{H}$ is a Hill angular momentum and $L_{c}$ is the clump angular momentum. The Hill angular momentum $L_{H}=M_{c}R_{H}^{2}\Omega_{K}$ where the Hill radius $R_{H}=a\big{(}M_{c}/(3M_{*})\big{)}^{\frac{1}{3}}$ for orbital semi-major axis $a$ and $\Omega_{K}=\sqrt{GM_{*}/a^{3}}$ the Keplerian angular rotation rate. The distribution of pebble cloud $l$ values in the streaming instability simulations has a mean value of 0.1, with the full range of values between 0.01 and 0.5 (see their Figure 4). The ratio of the clump angular momentum to a break up value associated with a maximum spin rate
$\frac{L_{c}}{L_{br}}=\frac{L_{c}}{L_{H}}\frac{L_{H}}{L_{br}}.$ (56)
We estimate $L_{br}\sim M_{c}R_{c}^{2}\Omega_{break-up}$ where $R_{c}$ is a final planetesimal radius after collapse and the maximum spin value is given in equation 3. We find that
$\frac{L_{H}}{L_{br}}\sim 1.5\left(\frac{\rho_{a}a^{3}}{M_{*}}\right)^{\frac{1}{6}}.$ (57)
This gives a Hill angular momentum to breakup ratio of
$\frac{L_{H}}{L_{br}}\sim 17\left(\frac{\rho_{a}}{1000~{}{\rm kg~{}m}^{-3}}\right)^{\frac{1}{6}}\left(\frac{M_{*}}{M_{\odot}}\right)^{-\frac{1}{6}}\left(\frac{a}{1~{}{\rm AU}}\right)^{\frac{1}{2}}.$ (58)
This ratio only depends upon orbital semi-major axis and the final planetesimal density. Inserting this into equation 56 and using the value of 0.1 for the ratio of cloud to Hill angular momentum measured in streaming instability simulations by Nesvorný et al. (2021), we estimate a cloud to maximum breakup angular momentum ratio of
$\frac{L_{c}}{L_{br}}\sim 2\left(\frac{L_{c}/L_{H}}{0.1}\right)\!\left(\frac{\rho_{a}}{1000~{}{\rm kg~{}m}^{-3}}\right)^{\frac{1}{6}}\!\!\left(\frac{M_{*}}{M_{\odot}}\right)^{-\frac{1}{6}}\!\!\left(\frac{a}{1~{}{\rm AU}}\right)^{\frac{1}{2}}.$ (59)
At 45 AU, $L_{c}/L_{br}\sim 13.4$, which is similar to the factor of $\sim$10 estimated by Nesvorný et al. (2021) in the Kuiper Belt, implying $\sim$ 93% of the cloud’s angular momentum needs to be removed to form a single planetesimal.
Equation 59 implies that the problem of angular momentum removal is worse in the outer solar system than in the inner solar system. In the outer solar system, due to the large fraction of angular momentum that must be lost, the angular momentum drain mechanism investigated in this paper would be incapable of facilitating collapse into a single planetesimal.
6 Summary
Planetesimals embedded in protostellar disks experience impacts from disk particles. If the planetesimal is rotating, escaping ejecta can reduce its spin angular momentum, as proposed in the context of asteroids by Dobrovolskis and Burns (1984). In this work we integrate the momentum of ejecta from a distribution of impacts to estimate the angular momentum carried away by ejecta. Our study differs in a number of ways from the work by Dobrovolskis and Burns (1984). Because impacts from particles in a disk would be at velocities of order 10-65 m/s rather than a few km/s, we take into account gravitational focusing. In this study we leverage scaling laws developed for crater ejecta distributions (Housen and Holsapple, 2011) that have been modified for oblique impacts so that they are sensitive to impact and azimuthal angle (Raducan et al., 2022; Quillen and Doran, 2024). In contrast Dobrovolskis and Burns (1984) neglected the azimuthal angular dependence of the ejecta distributions. Because the planetesimal has recently formed, we assume the direction from which the projectiles originate is related to the spin axis of the planetesimal and that all projectiles originate from the same direction. In contrast Dobrovolskis and Burns (1984) integrated over a distribution of projectiles that come from all directions. We consider impacts that cause accretion as well as those that cause erosion. In contrast, at the high impact velocities considered by Dobrovolskis and Burns (1984), the impacts only cause erosion.
We find that the angular momentum drain mechanism proposed by Dobrovolskis and Burns (1984) operates on planetesimals that are embedded within a circumstellar disk and are experiencing impacts originating from a headwind. We characterize the efficiency of the process by computing the ratio of the angular momentum lost per unit projectile mass and the planetesimal’s spin angular momentum. We find that the angular momentum drain mechanism is inefficient, with efficiency $\epsilon_{L}\ll 1$ and at most a few percent. Thus, if the impacting material is associated with a pebble cloud and contains angular momentum, the impact mediated angular momentum drain mechanism, would not significantly facilitate planetesimal collapse. With N-body simulations Nesvorný et al. (2021) found that the angular momentum of a pebble cloud prevents the collapse into a single planetesimal. Our study supports this result but with a different numerical approach. In our study the angular momentum transfer arises through ejecta scaling rather than dissipation associated with inelastic collisions.
We find the angular momentum transfer efficiency is most sensitive to headwind velocity and the ratio of planetesimal to projectile density. As planetesimal mass increases, the effects of gravitational focusing cause a decrease in the angular momentum transfer efficiency and an increase in the mass loss ratio due to the higher impact velocity caused by gravitational focusing. Assuming the projectiles originate from a disk filament, that might be present due to streaming instability with a velocity gradient capable of generating the computed pebble cloud, we find no meaningful reduction of the angular momentum transfer efficiency with regard to spinning down an embedded planetesimal. Future studies could take into account the angular momentum that is inherent in a population of projectiles.
The impact related mechanism of angular momentum loss, because it is inefficient, does not help resolve the issue of how excess angular momentum is shed in a collapsing pebble cloud. Additional sources of angular momentum drain may be necessary for the collapse of single-planetesimals. Pebble cloud collapse could lead preferentially to binary formation, as explored by Nesvorný et al. (2019, 2021). Alternatively, planetesimal formation itself could be inefficient with many pebble clouds failing to collapse, and only those that do collapse initially having low angular momentum.
Acknowledgements: This material is based upon work supported by NASA grant 80NSSC21K0143.
The python code used in this study is archived at http://xxxxarxiv.org will be updated upon posting to arxiv
Appendix A The Momentum and Angular Momentum Transfer
The momentum transfer efficiency, commonly called $\beta$, is a dimensionless parameter used to characterize deflection of an asteroid by an impact (e.g., Holsapple and Housen 2012; Feldhacker et al. 2017; Rivkin et al. 2021; Raducan et al. 2022). We recompute the momentum transfer parameter, taking into account the possibility of body rotation. We also compute the planetesimal’s angular momentum change caused by the impact.
We consider two sums, one for the initial total momentum of an impact in an inertial frame and one for the final total momentum,
Initial: ${\bf P}_{a}+m_{pj}{\bf u}_{pj}+M_{ej,esc}{\bf u}_{rot}$ (60)
Final: ${\bf P}_{a}+\Delta{\bf P}_{a}+M_{ej,esc}{\bf u}_{ej,esc}.$ (61)
Here ${\bf P}_{a}$ is the planetesimal’s initial momentum, but excluding that of the mass that will be ejected by the impact. The change of the planetesimal momentum is $\Delta{\bf P}_{a}$, $m_{pj}{\bf u}_{pj}$ is the projectile momentum at the moment of impact. The velocity ${\bf u}_{rot}$ is the rotational velocity of the surface at the site of impact, $M_{ej,esc}$ the total mass of escaping ejecta, and $M_{ej,esc}{\bf u}_{rot}$ is the initial momentum of the mass in escaping ejecta but prior to impact and still on the planetesimal surface and rotating with it. The escaping ejecta has momentum $M_{ej,esc}{\bf u}_{ej,esc}$ which we describe with a mean velocity ${\bf u}_{ej,esc}$ (in the inertial frame). After impact, we assume the projectile comes to rest on the surface but if part of it is ejected, that part could be incorporated into the escaping ejecta. Conservation of momentum implies that the initial and final momentum sums are equal, giving
$\Delta{\bf P}_{a}$ $=m_{pj}{\bf u}_{pj}+M_{ej,esc}({\bf u}_{rot}-{\bf u}_{ej,esc}).$ (62)
The velocity difference is equivalent to a velocity in the frame rotating with the surface (via equation 5). The momentum of escaping ejecta seen in the rotating frame
${\bf p}_{ej,esc}=M_{ej,esc}({\bf u}_{ej,esc}-{\bf u}_{rot})=M_{ej,esc}{\bf v}_{ej,esc}$ (63)
where ${\bf v}_{ej,esc}$ is the net velocity of escaping ejecta, as seen in the rotating frame. Thus the change in planetesimal momentum
$\Delta{\bf P}_{a}$ $=m_{pj}{\bf u}_{pj}-{\bf p}_{ej,esc}.$ (64)
The change in planetesimal velocity in the inertial frame, also called the deflection, would be
$\Delta{\bf V}_{a}=\frac{\Delta{\bf P}_{a}}{M_{a}}=\frac{m_{pj}{\bf u}_{pj}-{\bf p}_{ej,esc}}{M_{a}}.$ (65)
The momentum of escaping ejecta is often decomposed into two components, one normal to the surface and the other in the down-range or $\hat{\bf d}$ direction (Rivkin et al., 2021; Raducan et al., 2022). If the planetesimal is spherical, then the normal direction $\hat{\bf n}$ is the same in inertial and rotating frames and the momentum transfer efficiency parameter $\beta$ is consistent with
${\bf p}_{ej,esc}\cdot\hat{\bf n}=m_{pj}(1-\beta)(\hat{\bf n}\cdot{\bf u}_{pj}).$ (66)
Neglecting rotation, we recover expressions for $\beta$ and the deflection given by Feldhacker et al. (2017); Rivkin et al. (2021); Raducan et al. (2022).
We carry out a similar calculation but for the angular momentum. We find that the change in planetesimal angular momentum is
$\Delta{\bf L}_{a}$ $=m_{pj}{\bf r}\times{\bf u}_{pj}-{\bf r}\times{\bf p}_{ej,esc}$
$={\bf r}\times\Delta{\bf P}_{a}$ (67)
where $\bf r$ is the vector from the center of the planetesimal to the impact site. We confirm that a normal impact at longitude = 0, with projectile that sticks to the surface and gives no ejecta, would cause a change in angular momentum $\Delta{\bf L}_{a}=0$, as expected.
Assuming that the planetesimal is experiencing many collisions, if the distribution of projectiles has no angular momentum then the $m_{pj}{\bf r}\times{\bf u}_{pj}$ term in equation 67 will average to zero. We note that in equations 64 and 67, the relevant escaping ejecta momentum is computed in the frame rotating with the surface at the site of impact.
Up to this point, we have neglected gravitational focusing. Equations 60 and 61 describe the impact and ejecta curtain just prior to and after the impact. The projectile velocity ${\bf u}_{pj}$ refers to velocity of the projectile relative to the planetesimal at the moment of impact (in an inertial, not rotating frame) and is not equal to the velocity of the projectile at large distance from the planetesimal. As a consequence equation 66 describes the momentum transfer efficiency $\beta$ only for high velocity projectiles.
Equation 67 describes the difference between the angular momentum of the planetesimal projectile just before and after the impact. However, the gravitational force between planetesimal and projectile, which changes the relative momentum between projectile and planetesimal, conserves angular momentum. Thus equation 67 can be used to estimate the change in angular momentum by the impact and need not take into account the trajectories of ejecta escaping the planetesimal.
Appendix B The Hyperbolic Orbit Relevant for Gravitational Focusing
We consider a particle approaching a planetesimal. We neglect the gravitational force from other bodies and gas drag. The particle (or projectile) is not gravitationally bound to the planetesimal, so it is on a hyperbolic orbit. We describe a hyperbolic orbit of a zero mass projectile about a planetesimal of mass $M_{a}$ in terms of a semi-major axis $a$ and eccentricity $e$. The semi-major axis of the hyperbolic orbit is directly related to the orbital energy and can be expressed as
$a=-\frac{GM_{a}}{u^{2}_{hw}}.$ (68)
We solve for the orbital eccentricity $e$ in terms of impact parameter $b$, and the projectile relative velocity distant from the planetesimal $u_{hw}$:
$e^{2}=1+\frac{b^{2}u^{4}_{hw}}{(GM_{a})^{2}}.$ (69)
The distance from the center of mass of the planetesimal to the projectile is
$r(f)=\frac{a(1-e^{2})}{1+e\cos f}$ (70)
where $f$ is the true anomaly.
We solve for the radius where the orbit intersects the planetesimal surface, at a radius of $R_{a}$, giving the true anomaly of the impact site
$\cos f_{impact}=\frac{a(1-e^{2})}{R_{a}e}-\frac{1}{e}.$ (71)
The maximum true anomaly occurs distant from the planetesimal, when
$\cos f_{\infty}=-\frac{1}{e}.$ (72)
The angle between source wind direction (and the source of the projectile) and the site of impact, as measured from the planetesimal’s center of mass, is
$\Delta f=f_{\infty}-f_{impact}.$ (73)
This angle is used in section 2.8 to calculate the impact location, velocity and angle.